OpenStudy (anonymous):

Standardized Test Practice: Nine less than half n is equal to one plus the product -1/8th and n. Find the value of n. A. 24. Show work Please (:! Can somebody help because I have no idea how to do this!

OpenStudy (anonymous):

I can not figure this out.

OpenStudy (anonymous):

n-9=1+(-1/8n) ?? I know you are supposed to take the information and turn it into one equation so you can solve it but i got a little confused at the end.

OpenStudy (sasogeek):

ok so from the information we have, \(\large 9 \) less than \(\large \frac{1}{2}n \) is the same as reducing \(\large \frac{1}{2}n \) by \(\ 9 \) which is the same as \(\ large \frac{1}{2}n - 9 \). now we're told that this is equal to \(\large 1 \) plus the product of \(\large \frac{1}{8} \) and \(\large n \). product of that means \(\large \frac{1}{8} \times n\) which is the same as \(\large \frac{n}{8} \) . so back to the question, \(\ large \frac{n}{2} - 9 \) is equal to \(\large 1 + \frac{n}{8} \) \(\large \frac{n}{2} - 9 \)=\(\large 1 + \frac{n}{8} \) can you solve this equation on your own?